| <p>On 24/11/2018:</p>
<p>After the short test with 4.0MeV 7Li (see eLog #55) the beam was tuned for 4.0MeV 7Be 2+.</p>
<p>Run # 181124103943</p>
<ul>
<li>Run time: 43s</li>
<li>Target: CH2 XVI</li>
<li>Beam: 4.0MeV 7Be 2+</li>
<li>Comment: Far S2 iris was closed!</li>
</ul>
<p>Run # 181124104408</p>
<ul>
<li>Run time: 3971s</li>
<li>Pulser ON, as before</li>
<li>Target: CH2 XVI</li>
<li>Beam: 4.0MeV 7Be 2+</li>
</ul>
<p>After this run we checked the beam current:</p>
<p>FC4 = 10pA</p>
<p>Using 10mm aperture FCC = 4.5pA, using 3mm aperture FCC = 4.0pA</p>
<p>Attempted to increase beam current by adjusting bending magnet - optimum: 413.30mT</p>
<p>Run # 181124120558</p>
<ul>
<li>Run time 7930s</li>
<li>Pulser ON, as before</li>
<li>Target: CH2 XVI</li>
<li>Beam: 4.0MeV 7Be 2+</li>
</ul>
<p>See attached picture 1 for a sample spectrum from the S2 front strip. See attached picture 2 for the diode spectrum - the peaks from 7Be and 7Li are visible. Recall the diode has a 0.9um mylar protective foil in front of it.</p>
<p>Run # 181124141834</p>
<ul>
<li>Run time: 3362s</li>
<li>Pulser ON, as before</li>
<li>Target: CH2 XVI</li>
<li>Beam: 4.0MeV 7Be 2+</li>
</ul>
<p>Further attempts to increase beam current, ioniser increased from 24.0 -> 24.5A. FC4 = 15pA</p>
<p>Run # 181124154345</p>
<ul>
<li>Run time: 3609s</li>
<li>Pulser ON, as before</li>
<li>Target: CH2 XVI</li>
<li>Beam: 4.0MeV 7Be 2+</li>
</ul>
<p>Checked beam current: using 10mm aperture FCC = 4pA (time: 16:44)</p>
<p>Decreased oven temperature -> FC4 = 16pA</p>
<p>Run # 181124165221</p>
<ul>
<li>Run time: 4221s</li>
<li>Pulser ON, as before</li>
<li>Target: CH2 XVI</li>
<li>Beam: 4.0MeV 7Be 2+</li>
</ul>
<p>Collected background using carbon target:</p>
<p>Run # 181124180623</p>
<ul>
<li>Run time: 463s</li>
<li>Pulser ON, as before</li>
<li>Target: natC III</li>
<li>Beam: 4.0MeV 7Be 2+</li>
</ul>
<p>Beam current seemed unstable, run stopped, no actions taken. FCC ~ 5pA.</p>
<p>Run # 181124182304</p>
<ul>
<li>Run time: 3647s</li>
<li>Pulser ON, as before</li>
<li>Target: natC III</li>
<li>Beam: 4.0MeV 7Be 2+</li>
</ul>
<p>See attached picture 3 for a sample spectrum from the S2 front strip. The spectrum indicates hydrogen contamination is present on the carbon foil. See attached picture 4 for the diode spectrum.</p>
<p>After run: checked beam current. FC4 = 18pA. Using 10mm aperture FCC = 5.5pA.</p>
<p>Leaving for the night:</p>
<p>Far S2 leak I = 1.87uA. Far S2 and diode biases off.</p>
<p>19:25 pressures (mBar):</p>
<ul>
<li>TPU1 = 1.4e-6</li>
<li>TPU2 = 3.7e-7</li>
<li>TPD1 = 7.0e-7</li>
<li>TPD2 = 6.6e-7</li>
<li>CHUP = Sensor Off</li>
<li>CHDW = No Sensor</li>
</ul> |